A concrete bridge is built of 375-cm-long concrete slabs with an expansion joint between them. The slabs just touch on a 115°F day, the hottest day for which the bridge is designed Part A What is the gap between the slabs when the temperature is o°F? Express your answer with the appropriate units. ΔΙΞΙ Value Units Submit Request Answer
Answer #1

Please Check the value of alpha from your reference book, as it can be different in different books.

Relation between change in length and temperature is given by:

dL = L0*alpha*dT

L0 = original length of concrete slabs = 375 cm = 3.75 m

alpha = thermal expansion coefficient of concrete slabs = 12*10^-6 /C

dT = change in temperature = Tf - Ti

Tf = 0 F = -17.78 C

Ti = 115 F = 46.11 C

So, Using these values:

dL = 3.75*12*10^-6*(-17.78 - 46.11)

dL = -2.87*10^-3 m

dL = -2.87 mm

So gap at 0 F will be 2.87 mm

Please Upvote.